Taylor Smart    d.o.b. 1-15-19    wt.=24.5 lb.

Phenytoin 30mg/ml

Disp: 120ml

Sig: 4mg/kg/day in two divided doses.

How many milliliters of phenytoin would be administered to the child in each DOSE?
Answer #1

Weight=24.5 lbs( convert into kilograms multiply with 0 .453)

ie, 24.5 × 0.453=11.09 kg

Available dose=120 ml

In 1 ml= 30 mg

Recommended dose= 4 mg/kg/day in two divided doses

If 4 mg per kg ,then for 11.09 kg it is

11.09× 4 mg= 44.36 mg

For single dose = 44.36÷2= 22.18 mg

If 22.18 mg = X ml

Using proportion method

30 mg/1 ml= 22.8 mg/ X ml

Therefore, X =22.8 ÷30=0.739 = 0.74 ml

for one dose 0.74 ml should be administered.